OpenStudy (naudiapie):

Luther Vandekampt purchased a basketball from a store and gave the clerk a $20 bill and a $5 bill and received $1.12 in change. If the sales tax rate in his community is 6.5%, what was the selling price for the basketball before taxes?

OpenStudy (sweetburger):

So $23.88 is the selling price of the basketball including sales tax. So to find the actual price of the basketball we do this .065 *x = y Now take x+y = 23.88 x would end up being the actual price i honestly cant think of how to do this except for spamming different values for x and seeing which combo for x and y ends up being equivalent to 23.88

OpenStudy (sweetburger):

Theres probably an easier way to do it i just cant think of it.

OpenStudy (naudiapie):

$22.42 $22.88 $23.42 $23.88

OpenStudy (sweetburger):

Ahh that makes it much easier.

OpenStudy (naudiapie):

can you give me an easier equation then?

OpenStudy (shadowlegendx):

@sweetburger Should teach @Naudiapie how to solve it without relying on multiple choice answers

OpenStudy (sweetburger):

Well unfortunately the equation above will be the simplest to achieve the desired answer. Each of the answer choices a. $22.42 b. $22.88 c. $23.42 d. $23.88 represent values of x in the equation x*.065=y Ill try out answer choice a. to show you how to do it. 22.42*(.065) = y, so y = 1.46 x+y =23.88 22.42+1.46= (ill let you finish this)

OpenStudy (sweetburger):

@ShadowLegendX I would probably just try values of x that i would deem fit in my head and then move in towards a closer answer until i achieved the correct answer.

OpenStudy (naudiapie):

22.42+1.46= 23.88

OpenStudy (naudiapie):

okay i see :) thanks @sweetburger

OpenStudy (sweetburger):

Np glad to help